THE STUDY OF SMALL BIOMOLECULES USING CHIRPED-PULSE FOURIER TRANSFORM MI-CROWAVE (CP-FTMW) SPECTROMETER IN THE GAS PHASE^{a}

\underline{RYAN G. BIRD} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, Pa 15213; JUSTIN L. NEILL and BROOKS H. PATE, Department of Chemistry, University of Virginia, Charlottesville, Va 22904.

The microwave spectra of a large molecule is sensitive to both the structure of the molecule and as well as its motion along different vibrational coordinates. Chirped-pulse Fourier transform microwave (CP-FTMW) spectroscopy is an exciting new technique that makes possible the reading of the complete microwave spectrum of a gas phase sample using a single 1 s pulse. In this report, we will describe the recent introduction of a laser ablation nozzle for the study of small biomolecules using this technique. Potential applications of the technique to other samples also will be described.

aWork supported by NSF (CHE-0618740).